Search Mailing List Archives
[snf-lithography] Morning Litho Status 2/13/2014
Gary J Sosa
gsosa at stanford.edu
Thu Feb 13 08:04:38 PST 2014
Good morning All..
General Information/Lab Activities
Thursday, February 13, 2014
EE410 labs in progress . Please be advised that the EE410 classes have priority on the tools. Please observe and respect the tool reservations for these lab sessions.
Mario - Out today(2/13)
Gary - Out @ 3:00PM Today
Scheduled lamp change today
Does not work as standard aligner. Suspect software or configuration
No actions planned at this time
Up for 4” and 6” wafer processing
No Argon flow
Replaced TC gauge and meter with a suitable replacement because original parts are obsolete. Tested N2 and Argon functions-OK. System needs a process test.
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the snf-lithography